OpenStudy (anonymous):

Solve for t: xf = xi + at^2

OpenStudy (anonymous):

at^2=xf-xi t^2=(xf-xi)/a t=sqrt ((xf-xi)/a)

OpenStudy (raden):

dont forget give sign +- in front sqrt @stgreen

OpenStudy (anonymous):

dude time can never be negative.....common sense

OpenStudy (anonymous):

^ @RadEn

OpenStudy (raden):

why not, if u square it u will back to the original problem :p

OpenStudy (anonymous):

lol have you ever talked like "-4 mins"??

OpenStudy (raden):

and do u agree like this : x^-4=0 so, x=+-2, it means x satisfies for x=-2 or x=2

OpenStudy (raden):

x^2-4=0 i meant :)

OpenStudy (anonymous):

but TIME is nver negative na.....clocks don't go back....

OpenStudy (anonymous):

Thanks for making it easy. I was stuck on the 2nd step for a minute there